nono:

idk to slothfull A florist has an order to make flower bouquets. The order calls for the use of 45 roses and 63 carnations. Each bouquet made will need to be the same. If all of the flowers are to be used, there can be at most bouquets made from the flowers requested in the order.

sillybilly123:

This is a Binomial Distribution.

sillybilly123:

Perhaps!

Vocaloid:

Since they must use ALL of the flowers I would find the GCF between the two numbers 63 = 7*9 45 = 9*5

Vocaloid:

The gcf is the number of bouquets and the leftover factors are the number of flowers of each corresponding type

Vocaloid:

Closed due to answered. Please re-open a new question if you still need assistance.